### v2

# Masked Hard-Attention Transformers and Boolean RASP Recognize Exactly the Star-Free Languages

Abstract.
We consider transformer encoders with hard attention (in which all attention is focused on exactly one position) and strict future masking (in which each position only attends to positions strictly to its left), and prove that the class of languages recognized by these networks is exactly the star-free languages. Adding position embeddings increases the class of recognized languages to other well-studied classes. A key technique in these proofs is Boolean RASP, a variant of RASP that is restricted to Boolean values. Via the star-free languages, we relate transformers to first-order logic, temporal logic, and algebraic automata theory.
conference: Thirty-Ninth Annual ACM/IEEE Symposium on Logic in Computer Science (LICS); 8–12 July 2024; Tallinn, Estonia
1. Introduction
[…]
Figure 1. Summary of previous results about star-free languages, and new results about masked hard-attention transformers and B-RASP, without position embeddings.
Significant progress has been made in the last few years on characterizing the expressivity of transformers (Vaswani et al., 2017) in terms of well-understood classes of formal languages. Results have been obtained for a wide range of variants of transformers, and nearly all take the form of either upper bounds (transformers recognize only languages in class $C$) or lower bounds (transformers recognize all languages in class $C$). In this paper, we establish exact characterizations of transformer encoders with hard attention (in which all attention is focused on exactly one position), and strict future masking (in which each position only attends to positions strictly to its left).
Without position embeddings, these transformers recognize exactly the class of star-free regular languages. Figure 1 summarizes previous results about star-free languages and the results added in this paper. The class of star-free regular languages is an especially robust one, with multiple diverse characterizations. Thus, our results connect transformers not only with the theory of regular expressions, but also first-order logic, temporal logic, and algebraic automata theory.
Adding position embeddings increases the class of recognized languages to other well-studied classes (Section 8.1). With sinusoidal position embeddings, masked hard-attention transformers are equivalent to $\textbf{FO}[\mathord{<},\text{MOD}]$ (first-order logic with modular predicates), or, equivalently, they recognize the regular languages in $\textbf{AC}^{0}$. With arbitrary position embeddings, they are equivalent to $\textbf{FO}[\mathord{<},\text{Mon}]$ (first-order logic with arbitrary monadic predicates).
A key technique in these proofs is the use of B-RASP, a variant of RASP (Weiss et al., 2021) that is restricted to Boolean values. Like RASP, it is a small programming language that compiles into transformers, for the purpose of more easily constructing transformers by hand. B-RASP compiles to masked, Boolean-valued transformers, and unlike with RASP, the correspondence is exact, in the sense that a masked Boolean transformer can be decompiled back to a B-RASP program. We use B-RASP as a convenient intermediate language between transformers and other characterizations of star-free languages.
2. Background
2.1. Transformer variants
The original transformer (Vaswani et al., 2017), designed for machine translation, had both an encoder and a decoder. Pérez et al. (2021) show that transformer encoder–decoders, where the decoder runs for a polynomial number of steps before accepting or rejecting a string, recognize all of P. Merrill and Sabharwal (2024) prove another version of the same result, and they observe that all such transformers are in P. This result, contemporaneous with ours, is the only other exact characterization of any transformer variant that we are aware of.
Encoder-only models like BERT (Devlin et al., 2019) have been widely used in practice, as have decoder-only models like GPT (Brown et al., 2020). Most previous work on transformer expressivity have studied encoder-only models (e.g. Hahn, 2020), and the present paper does as well.
The transformer encoders studied here use hard attention, also known as unique hard attention, in which an attention head focuses all attention on the position with the highest score, with ties broken to the left or right. Hard attention has been incorporated into transformer models in previous works on interpretability (Kinley, 2020) and efficiency (Gupta et al., 2021). Hao et al. (2022) show that hard-attention transformers with arbitrary position embeddings have an upper bound of $\textbf{AC}^{0}$, and, in work contemporaneous with ours, Barceló et al. (2024) show that they have a lower bound of $\textbf{FO}[\mathord{<},\text{Mon}]$, which is first-order logic with all possible monadic numerical predicates. They leave open the question of whether these transformers are equivalent to $\textbf{FO}[\mathord{<},\text{Mon}]$—a question which, with suitable adjustments, we answer here in the affirmative.
In this paper, we use future masking, in which every position may only attend to positions to the left. This kind of masking is commonly used in decoder-only models and has been studied in encoder-only models as well (Bhattamishra et al., 2020). We also consider past masking (Yao et al., 2021). Our attention masking is strict, that is, each position may not attend to itself. As far as we know, this is new, and, interestingly, in the setting of hard-attention transformers, strict future/past masking is more powerful than non-strict (Section 5.5).
2.2. Notation
Let $\Sigma$ be a finite alphabet, and let $w=w_{1}\cdots w_{n}$ be an input string of length $n$, where each $w_{i}\in\Sigma$. We often assume a distinct end-of-string symbol EOS not in $\Sigma$; the set $\Sigma\cup\{\texttt{EOS}\}$ is denoted $\Sigma_{\texttt{EOS}}$. We write $[n]$ for the set $\{1,\ldots,n\}$.
2.3. Star-free languages
The star-free regular languages can be defined as the closure of $\emptyset$, $\{\epsilon\}$, and $\{a\}$ for each $a\in\Sigma$, under the operations of union ($A\cup B$), concatenation ($AB$), and complementation ($A^{\textsf{C}}$). For example:
  • $\Sigma^{*}$ is star-free because $\Sigma^{*}=\emptyset^{\textsf{C}}$.
  • $a^{*}$ is star-free because $a^{*}=(\Sigma^{*}a^{\textsf{C}}\Sigma^{*})^{\textsf{C}}$.
  • $(ab)^{*}$ is star-free because $(ab)^{*}=(b\Sigma^{*}\cup\Sigma^{*}a\cup\Sigma^{*}aa\Sigma^{*}\cup\Sigma^{*}bb\Sigma^{*})^{\textsf{C}}$.
An example of a regular language that is not star-free is $(aa)^{*}$.
The class of star-free languages has a number of alternative characterizations, which we will introduce later in the paper.
Bhattamishra et al. (2020) investigate what regular languages transformers (specifically, transformer encoders with future masking) can recognize or learn to recognize. They find that, in practice, transformers can only learn star-free languages, and more specifically, only languages with dot-depth 1 (Cohen and Brzozowski, 1971). For example, $L_{1,2}$, the Dyck language with 1 pair of parentheses, limited to depth 2, has dot-depth 2, and Bhattamishra et al. (2020) argue that transformers cannot recognize it. We discuss this example in detail in Section 3.2.1.
3. B-RASP
RASP (Weiss et al., 2021) is a programming language intended to help programmers “think like transformers.” It has the same basic operations as transformers, but it is easier to compose these operations in RASP than by writing transformers by hand. However, RASP targets a variant of transformers that is more general than standard transformers, and there is no decompiler from transformers to RASP.
In this section, we define a version of RASP  restricted to Boolean values, which we call Boolean RASP or B-RASP. It compiles into Boolean-valued transformers, and by contrast with RASP, we also give the inverse transformation, from masked Boolean transformers to B-RASP.
3.1. Definition
To decide whether a B-RASP program accepts an input string $w=w_{1}\cdots w_{n}\in\Sigma^{*}$, the string $w\cdot\texttt{EOS}$ is presented as input to the program. The EOS serves three purposes: (i) it provides a position where the output of a program is found (similar to CLS in a BERT-based classifier); (ii) it serves as a default position when no positions can be attended to; (iii) it simplifies the case of an empty input string, which is presented to the program as EOS.
There is one type of data, a Boolean vector, which is a vector of Boolean values indexed by $i\in[n+1]$. Initially, there is one Boolean vector $Q_{a}$ for each $a\in\Sigma_{\texttt{EOS}}$, where $Q_{a}(i)=1$ iff $w_{i}=a$.
A B-RASP program is a sequence of definitions of new Boolean vectors. The Boolean vectors may have descriptive names, and names may be reused; but to streamline the definition of B-RASP and proofs about B-RASP, we assume that all the Boolean vectors are numbered consecutively, so that $P_{1},\ldots,P_{|\Sigma|+1}$ are the initial Boolean vectors $Q_{a}$ for $a\in\Sigma_{\texttt{EOS}}$, and the Boolean vectors defined by the program are numbered from $P_{|\Sigma|+2}$ without repetition.
Assume that Boolean vectors $P_{1},\ldots,P_{k}$ have been defined. Then $P_{k+1}$ is defined using one of the following forms.
Position-wise operations. A Boolean vector $P_{k+1}(i)$ can be a Boolean combination of zero or more of $\{P_{1}(i),\ldots,P_{k}(i)\}$.
Attention operations. An attention operation can be either of
$P_{k+1}(i)$ $={\operatorname{{\mathbf{\blacktriangleleft}}}}\left[M(i,j),S(i,j)\right]\;V(i,j)$
$P_{k+1}(i)$ $={\operatorname{{\mathbf{\blacktriangleright}}}}\left[M(i,j),S(i,j)\right]\;V(i,j)$
where:
  • $M(i,j)$ is a mask predicate, one of
    • no masking: $M(i,j)=1$
    • future masking: $M(i,j)=(j<i)$
    • past masking: $M(i,j)=(j>i)$.
  • $S(i,j)$ is a score predicate, which is a Boolean combination of zero or more atomic formulas from $\{P_{1}(i),\ldots,P_{k}(i)\}\cup\{P_{1}(j),\ldots,P_{k}(j)\}$.
  • $V(i,j)$ is a value predicate, which has the same form as the score predicate.
[…]
If $P$ is a Boolean vector computed by program $p$, we write $w\models P(i)$ just in case $P(i)=1$ when $p$ is run on input string $w$.
To make a B-RASP program $p$ recognize a language, one Boolean vector $Y$ is designated as the output vector and position $(n+1)$ is designated the output position. Then, the input string $w$ is accepted iff $w\models Y(n+1)$.
To enable B-RASP to compute length-preserving sequence-to-sequence functions from $\Sigma^{*}$ to $\Gamma^{*}$, we designate a collection of output Boolean vectors $Y_{\gamma}$ indexed by the symbols $\gamma\in\Gamma$, and consider the final values of the Boolean vectors $Y_{\gamma}(i)$ for $i\in[n+1]$ to be one-hot encodings of the sequence of output symbols.
3.2. Examples
3.2.1. Dyck-$1$ of depth 22
The first example we consider is the language Dyck-$1$ of depth $2$. This is the regular language $L_{1,2}$ accepted by the DFA in Figure 2, where $\ell$ and $r$ are left and right brackets. Our construction is very similar to that of Yao et al. (2021).
$\ell\ellrrrr$
Figure 2. DFA recognizing $L_{1,2}$, Dyck-11 of depth $2$. The left bracket is $\ell$ and the right bracket is $r$.
Consider the input string $\ell\ell rr\ell\ell r\ell rr$, which should be accepted. The basic idea of the construction is to replace brackets that are immediately matched by 0, which yields $\ell 00r\ell 0000r$, and then to make sure the remaining brackets are properly matched. We describe the B-RASP program for this problem below; the resulting Boolean vectors are shown in Figure 3.
input $\ell$ $\ell$ $r$ $r$ $\ell$ $\ell$ $r$ $\ell$ $r$ $r$ EOS
$Q_{\texttt{EOS}}$ 0 0 0 0 0 0 0 0 0 0 1
$Q_{\ell}$ 1 1 0 0 1 1 0 1 0 0 0
$Q_{r}$ 0 0 1 1 0 0 1 0 1 1 0
$P_{\ell}$ 0 1 1 0 0 1 1 0 1 0 0
$S_{r}$ 0 1 1 0 0 1 0 1 1 0 0
II 0 1 1 0 0 1 1 1 1 0 0
$V_{\ell}$ 0 0 0 0 0 0 0 0 0 0 0
$V_{r}$ 0 0 0 0 0 0 0 0 0 0 0
$Y$ 1 1 1 1 1 1 1 1 1 1 1
Figure 3. Boolean vectors for membership of string $\ell\ell rr\ell\ell r\ell rr$ in $L_{1,2}$.
We first construct the Boolean vector $P_{\ell}$ that indicates whether the symbol immediately to the left is $\ell$. This is done with an attention operation:
$P_{\ell}(i)={\operatorname{{\mathbf{\blacktriangleright}}}}\left[j<i,1\right]\;Q_{\ell}(j).$
[…]
$S_{r}(i)={\operatorname{{\mathbf{\blacktriangleleft}}}}\left[j>i,1\right]\;Q_{r}(j).$
The Boolean vector $I$ indicates whether position $i$ is in a consecutive pair $\ell r$, that is, whether it is immediately matched.
$I=(Q_{\ell}\wedge S_{r})\vee(Q_{r}\wedge P_{\ell}).$
The Boolean vector $V_{\ell}$ registers a violation for position $i$ if it has symbol $\ell$, is not immediately matched, and the next following not immediately matched symbol is not $r$:
$V_{\ell}(i)$ $={\operatorname{{\mathbf{\blacktriangleleft}}}}\left[j>i,\neg I(j)\right]\;(Q_{\ell}(i)\wedge\neg I(i)\wedge\neg Q_{r}(j)).$
The Boolean vector $V_{r}$ registers a violation for position $i$ if it has symbol $r$, is not immediately matched, and the preceding not immediately matched symbol is not $\ell$:
$V_{r}(i)$ $={\operatorname{{\mathbf{\blacktriangleright}}}}\left[j<i,\neg I(j)\right]\;(Q_{r}(i)\wedge\neg I(i)\wedge\neg Q_{\ell}(j))$
The output Boolean vector $Y$ is computed using an attention operation:
$Y(i)$ $={\operatorname{{\mathbf{\blacktriangleleft}}}}\left[1,V_{\ell}(j)\vee V_{r}(j)\right]\;\neg(V_{\ell}(j)\vee V_{r}(j))$
Position $i$ attends to the leftmost position $j$ that registers a violation (if any). If that position registers a violation, then the output should be 0, otherwise the output should be $1$.
input $\ell$ $r$ $r$ $\ell$ $\ell$ $\ell$ $r$ $r$ $r$ $\ell$ EOS
$Q_{\texttt{EOS}}$ 0 0 0 0 0 0 0 0 0 0 1
$Q_{\ell}$ 1 0 0 1 1 1 0 0 0 1 0
$Q_{r}$ 0 1 1 0 0 0 1 1 1 0 0
$P_{\ell}$ 0 1 0 0 1 1 1 0 0 0 1
$S_{r}$ 1 1 0 0 0 1 1 1 0 0 0
II 1 1 0 0 0 1 1 0 0 0 0
$V_{\ell}$ 0 0 0 1 0 0 0 0 0 1 0
$V_{r}$ 0 0 1 0 0 0 0 0 1 0 0
$Y$ 0 0 0 0 0 0 0 0 0 0 0
Figure 4. Boolean vectors for non-membership of string $\ell rr\ell\ell\ell rrr\ell$ in $L_{1,2}$.
Boolean vectors for deciding non-membership of $\ell rr\ell\ell\ell rrr\ell$ in $L_{1,2}$ are shown in Figure 4. It is straightforward to generalize this technique to recognize Dyck-$k$ of depth $D$ in B-RASP. [^1] Because we prove in Theorem 7.1 that B-RASP programs can be simulated by masked hard-attention transformers, this result contradicts the claim in Theorem 4.3 (= Theorem C.1) of the paper by Yao et al. (2021); according to a cognizant co-author of that paper, Lemma C.2 in that paper is not true (Peng, 2023).
3.2.2. A retrieval task
We consider a variation of the retrieval task studied by Friedman et al. (2023). Inputs are strings over the alphabet $\{a,b,c,1,2,3\}$, where letters and numbers alternate, starting with a letter and ending with a number, for example $w=a3b2b1a2c1a1c3$. The output alphabet adds the symbol ??. The desired output sequence copies the letters, and for a number at position $i$, if $\sigma$ is the letter at position i1i-1, then the most recent previous occurrence of $\sigma$ is found, say at position $j$, and the number at position $j+1$ is output. If there is no previous occurrence of $\sigma$, then ?? is output instead. For the given input $w$, the output should be $y=a?b?b2a3c?a2c1$.
The Boolean vector $P_{a}$ determines whether there is an $a$ in the preceding position.
$P_{a}(i)={\operatorname{{\mathbf{\blacktriangleright}}}}\left[j<i,1\right]\;Q_{a}(j).$
The Boolean vector $P_{b}$ (resp., $P_{c}$) indicating whether there is a $b$ (resp., $c$) in the preceding position is computed similarly.
For $k\in\{1,2,3\}$, the Boolean vector $Y_{k}$ indicates whether the output symbol is $k$. If position $i$ is preceded by a letter, it attends to the most recent preceding position $j$ preceded by the same letter (if any), and sets $Y_{k}(i)$ to the value of $Q_{k}(j)$ for that $j$:
$Y_{k}(i)$ $={\operatorname{{\mathbf{\blacktriangleright}}}}\left[j<i,\;\bigvee_{\mathclap{\sigma\in\{a,b,c\}}}P_{\sigma}(i)\land P_{\sigma}(j)\right]\;Q_{k}(j).$
For $\sigma\in\{a,b,c,\texttt{EOS}\}$, the output Boolean vector $Y_{\sigma}$ is just $Q_{\sigma}$. The output Boolean vector $Y_{?}$ is true if the position is preceded by a letter but no number was assigned:
$Y_{?}=(P_{a}\vee P_{b}\vee P_{c})\wedge\neg(Y_{1}\vee Y_{2}\vee Y_{3}).$
Selected vectors in the computation for the example are in Figure 5.
input $a$ 33 $b$ 22 $b$ 11 $a$ 22 $c$ 11 $a$ 11 $c$ 33 EOS
$P_{a}$ 0 1 0 0 0 0 0 1 0 0 0 1 0 0 0
$P_{b}$ 0 0 0 1 0 1 0 0 0 0 0 0 0 0 0
$P_{c}$ 0 0 0 0 0 0 0 0 0 1 0 0 0 1 0
$Y_{1}$ 0 0 0 0 0 0 0 0 0 0 0 0 0 1 0
$Y_{2}$ 0 0 0 0 0 1 0 0 0 0 0 1 0 0 0
$Y_{3}$ 0 0 0 0 0 0 0 1 0 0 0 0 0 0 0
$Y_{?}$ 0 1 0 1 0 0 0 0 0 1 0 0 0 0 0
output $a$ ?? $b$ ?? $b$ 22 $a$ 33 $c$ ?? $a$ 22 $c$ 11 EOS
Figure 5. Selected Boolean vectors for retrieval task on input $a3b2b1a2c1a1c3$.
4. B-RASP recognizes the star-free languages
We prove that B-RASP recognizes exactly the star-free languages. One direction of the proof (B-RASP to star-free) is via first-order logic, which defines exactly the star-free languages. We provide two proofs of the other direction (star-free to B-RASP). In this section, we prove it via linear temporal logic, which also defines exactly the star-free languages. Later, in Section 9, we will prove it via the Krohn-Rhodes decomposition for counter-free automata.
4.1. First-order logic
We consider languages over $\Sigma$ that are definable in first order logic ($\textbf{FO}[\mathord{<}]$). We have the usual connectives $\land$, $\lor$, and $\neg$, and quantifiers $\forall$ and $\exists$, and we write $p\rightarrow q$ for $\neg p\lor q$. We have predicates $Q_{a}(i)$ for each $a\in\Sigma$, and we have binary predicates == and <<.
Let $w=w_{1}\cdots w_{n}$ be an input string, Under the intended interpretation, variables range over $[n]$ and stand for positions of $w$.
If $\phi$ is a formula of $\textbf{FO}[\mathord{<}]$, and $\xi$ is an assignment of positions in $[n]$ to free variables of $\phi$, we write $w,\xi\models\phi$ if:
  • $\phi=Q_{a}(x)$ and $w_{\xi(x)}=a$
  • $\phi=(x<y)$ and $\xi(x)<\xi(y)$
  • $\phi=(x=y)$ and $\xi(x)=\xi(y)$
  • $\phi=(\phi_{1}\land\phi_{2})$ and $w,\xi\models\phi_{1}$ and $w,\xi\models\phi_{2}$
  • $\phi=(\phi_{1}\lor\phi_{2})$ and either $w,\xi\models\phi_{1}$ or $w,\xi\models\phi_{2}$
  • $\phi=\neg\phi_{1}$ and $w,\xi\not\models\phi_{1}$
  • $\phi=(\forall x)\phi_{1}$ and for all $i\in[n]$, $w,\xi[x\mapsto i]\models\phi_{1}$.
If $\phi$ has no free variables, we write $w\models\phi$ for $w,\emptyset\models\phi$, and if $\phi$ has just one free variable, we write $w\models\phi(i)$ for $w,\{x\mapsto i\}\models\phi$.
For example, we can define predicates
$\text{First}(i)$ $\equiv\neg(\exists j)(j<i)$
$\text{Succ}(i,j)$ $\equiv(j>i\land\neg(\exists k)(i<k\land k<j)).$
Then, as an example of a language definable in $\textbf{FO}[\mathord{<}]$, $(ab)^{*}$ is defined by the sentence
$(\forall i)(\begin{aligned} &(\text{First}(i)\rightarrow Q_{a}(i))\land{}\\ &(Q_{a}(i)\rightarrow(\exists j)(\text{Succ}(i,j)\land Q_{b}(j)))\land{}\\ &(Q_{b}(i)\rightarrow\neg(\exists j)(\text{Succ}(i,j)\land Q_{b}(j)))).\end{aligned}$
A famous result is that the formal languages definable in $\textbf{FO}[\mathord{<}]$ are precisely the star-free regular languages (McNaughton and Papert, 1971).
4.2. B-RASP to $\textbf{FO}[\mathord{<}]$
A B-RASP program $p$ is representable in $\textbf{FO}[\mathord{<}]$ if for each Boolean vector $P_{k}(i)$ in $p$, there exists a first-order definable predicate $P_{k}^{\prime}(i)$ such that for any input $w\in\Sigma^{*}$ of length $n$, for all $i\in[n+1]$, we have that $w\cdot\texttt{EOS}\models P_{k}^{\prime}(i)$ iff $w\models P_{k}(i)$.
Theorem 4.1.
Any program $p$ in B-RASP is representable in $\textbf{FO}[\mathord{<}]$.
Proof.
We prove that the first $k$ Boolean vectors of $p$ are representable in $\textbf{FO}[\mathord{<}]$, by induction on $k$. The base case $k=0$ is trivial. Assume that the Boolean vectors $P_{1},\ldots,P_{k}$ are representable in $\textbf{FO}[\mathord{<}]$. We want to show that $P_{k+1}$ is representable as well.
If $P_{k+1}$ is a built-in Boolean vector $Q_{a}$, it is represented by the corresponding first-order predicate $Q_{a}$.
If $P_{k+1}$ is a position-wise operation
$P_{k+1}(i)=f(P_{1}(i),\ldots,P_{k}(i))$
where $f$ is some Boolean function, then let
$P_{k+1}^{\prime}(i)=f(P_{1}^{\prime}(i),\ldots,P_{k}^{\prime}(i)).$
If $P_{k+1}$ is an attention operation
$P_{k+1}(i)={\operatorname{{\mathbf{\blacktriangleleft}}}}\left[M(i,j),S(i,j)\right]\;V(i,j)$
then we define a first-order predicate $P_{k+1}^{\prime}$ representing $P_{k+1}$ as follows. Let $S^{\prime}(i,j)$ and $V^{\prime}(i,j)$ be the first order predicates obtained from $S(i,j)$ and $V(i,j)$ by replacing each Boolean vector $P_{m}$ by the predicate $P^{\prime}_{m}$ that represents it. Then:
$P^{\prime}_{k+1}(i)$ $=\begin{aligned} &\neg(\exists j)M(i,j)\land(\exists z)(Q_{\texttt{EOS}}(z)\land V^{\prime}(i,z))\\ &{}\lor(\exists j)(M(i,j)\land\neg(\exists\ell)((\ell<j)\land M(i,\ell)\\ &\qquad\land(\neg S^{\prime}(i,j)\lor S^{\prime}(i,\ell)))\land V^{\prime}(i,j)).\end{aligned}$
That is, if there is no $j$ such that $M(i,j)$ is true, then the value of $P_{k+1}^{\prime}(i)$ is $V^{\prime}(i,n+1)$. Otherwise, among $j$ such that $M(i,j)=1$ there is a unique least $j$ that takes on the maximum value of $S(i,j)$, call it $j_{i}$, and the value of $P_{k+1}^{\prime}(i)$ is $V^{\prime}(i,j_{i})$.
If the operation is ${\operatorname{{\mathbf{\blacktriangleright}}}}\left[M(i,j),S(i,j)\right]\;V(i,j)$, then $(\ell<j)$ is replaced by $(j>\ell)$. ∎
Corollary 4.2.
Let $L\subseteq\Sigma^{*}$ be recognized by the B-RASP program $p$. Then $L$ is star-free.
Proof.
Suppose the output Boolean vector for $p$ is $Y$. Let $Y^{\prime}$ be the first-order logic formula that represents $Y$, and let $L^{\prime}$ denote the language over $\Sigma_{\texttt{EOS}}$ defined by $Y^{\prime}$. For every $w\in\Sigma^{*}$, $w\in L$ if and only if $w\cdot\texttt{EOS}\in L^{\prime}$. Thus, $L^{\prime}\cap(\Sigma^{*}\cdot\texttt{EOS})=L\cdot\texttt{EOS}$ and this language is star-free because $L^{\prime}$ and $(\Sigma^{*}\cdot\texttt{EOS})$ are star-free. Because $L\cdot\texttt{EOS}$ is star-free, so is $L$. ∎
4.3. Linear temporal logic to B-RASP
For the other direction, we use the fact that the star-free languages can be expressed by formulas in linear temporal logic (LTL), which defines the same languages as $\textbf{FO}[\mathord{<}]$ (Kamp, 1968).
In LTL, every formula implicitly depends on a single “time” (or position) variable. The atomic formulas are $Q_{a}$ for every $a\in\Sigma$, and we have the usual connectives $\land$, $\lor$, and $\neg$, as well as operators $\mathrel{\textbf{since}}$ and $\mathrel{\textbf{until}}$. For any input string $w=w_{1}\cdots w_{n}$ and position $i\in[n]$, we write $w,i\models\phi$ if:
  • $\phi=Q_{a}$ and $w_{i}=a$.
  • $\phi=(\phi_{1}\land\phi_{2})$ and $w,i\models\phi_{1}$ and $w,i\models\phi_{2}$.
  • $\phi=(\phi_{1}\lor\phi_{2})$ and either $w,i\models\phi_{1}$ or $w,i\models\phi_{2}$.
  • $\phi=\neg\phi_{1}$ and $w,i\not\models\phi_{1}$.
  • $\phi=\phi_{1}\mathrel{\textbf{since}}\phi_{2}$, and there is a $j<i$ such that $w,j\models\phi_{2}$ and for all $k$ such that $j<k<i$, $w,k\models\phi_{1}$.
  • $\phi=\phi_{1}\mathrel{\textbf{until}}\phi_{2}$, and there is a $j>i$ such that $w,j\models\phi_{2}$ and for all $k$ such that $i<k<j$, $w,k\models\phi_{1}$.
To use a formula $\phi$ of LTL to define a language over $\Sigma$, for a $w\in\Sigma^{*}$ of length $n$ we supply $w\cdot\texttt{EOS}$ as input and designate the last position as the output position, so that $w\in\mathcal{L}(\phi)$ if and only if $w\cdot\texttt{EOS},n+1\models\phi$. In this setting, the logic with just $\mathrel{\textbf{since}}$, known as past temporal logic, is as expressive as the logic with both operators (Gabbay et al., 1980; Maler and Pnueli, 1990), and recognizes all the star-free languages.
Theorem 4.3.
If a language can be defined by a formula of linear temporal logic, then it can be defined by a B-RASP program.
Proof.
By induction on the structure of the formula $\phi$.
If $\phi=Q_{a}$: Assume that the B-RASP program already defines $P_{1},\ldots,P_{k}$. Then add instruction
$P_{k+1}(i)=Q_{a}(i).$
If $\phi=\neg\phi_{1}$: By the induction hypothesis, convert $\phi_{1}$ to B-RASP instructions ending with a definition of $P_{k_{1}}$. Then add instruction
$P_{k_{1}+1}(i)=\neg P_{k_{1}}(i).$
If $\phi=\phi_{1}\land\phi_{2}$: By the induction hypothesis, convert $\phi_{1}$ to B-RASP instructions ending with a definition of $P_{k_{1}}$, then convert $\phi_{2}$ to B-RASP instructions ending with a definition of $P_{k_{2}}$. Then add instruction
$P_{k_{2}+1}(i)=P_{k_{1}}(i)\land P_{k_{2}}(i).$
If $\phi=\phi_{1}\lor\phi_{2}$: Similar, but add
$P_{k_{2}+1}(i)=P_{k_{1}}(i)\lor P_{k_{2}}(i).$
If $\phi=\phi_{1}\mathrel{\textbf{since}}\phi_{2}$: Similar, but add
$P_{k_{2}+1}(i)=\begin{aligned} {\operatorname{{\mathbf{\blacktriangleright}}}}&\left[j<i,\neg P_{k_{1}}(j)\lor P_{k_{2}}(j)\right]\\ &(P_{k_{2}}(j)\land\neg Q_{\texttt{EOS}}(j)).\end{aligned}$
If $\phi=\phi_{1}\mathrel{\textbf{until}}\phi_{2}$: Similar, but add
$P_{k_{2}+1}(i)=\begin{aligned} {\operatorname{{\mathbf{\blacktriangleleft}}}}&\left[j>i,\neg P_{k_{1}}(j)\lor P_{k_{2}}(j)\right]\\ &(P_{k_{2}}(j)\land\neg Q_{\texttt{EOS}}(j)).\end{aligned}$
The output of the B-RASP program is the Boolean vector constructed to represent $\phi$. ∎
Corollary 4.4.
If $L$ is a star-free language over $\Sigma$ then there exists a B-RASP program to recognize $L$.
5. B-RASP Variants
We consider some variants of B-RASP and show how all but one are equally expressive to our definition.
5.1. Asymmetric attention
[…]
5.2. Leftmost/rightmost true
[…]
$P^{\prime}_{k+1}(i)$ $=\begin{aligned} &\neg(\exists j)(M(i,j)\land S^{\prime}(i,j))\land(\exists z)(Q_{\texttt{EOS}}(z)\land V^{\prime}(i,z))\\ &{}\lor(\exists j)(M(i,j)\land S^{\prime}(i,j)\\ &\qquad\land\neg(\exists\ell)((\ell<j)\land M(i,\ell)\land S^{\prime}(i,\ell))\land V^{\prime}(i,j)).\end{aligned}$
[…]
The above simulation of LTL $\mathrel{\textbf{since}}$ and $\mathrel{\textbf{until}}$ (Section 4.3) also works for this variant.
5.3. Default value
Another possibility is a default value instead of a default position. In this case, an attention operation looks like
${\operatorname{{\mathbf{\blacktriangleleft}}}}\left[M(i,j),S(i,j)\right]\;V(i,j)\mathrel{\textbf{else}}V_{\text{default}}(i)$
If all positions are masked out, the attention operation returns $V_{\text{default}}(i)$.
[…]
$P^{\prime}_{k+1}(i)$ $=\begin{aligned} &\neg(\exists j)(M(i,j)\land S^{\prime}(i,j))\land V_{\text{default}}(i)\\ &{}\lor(\exists j)(M(i,j)\land S^{\prime}(i,j)\\ &\qquad\land\neg(\exists\ell)((\ell<j)\land M(i,\ell)\land S^{\prime}(i,\ell))\land V^{\prime}(i,j)).\end{aligned}$
And we can simulate LTL $\mathrel{\textbf{until}}$:
$P_{k_{2}+1}(i)=\begin{aligned} {\operatorname{{\mathbf{\blacktriangleleft}}}}\left[j>i,\neg P_{k_{1}}(j)\lor P_{k_{2}}(j)\right]\;P_{k_{2}}(j)\mathrel{\textbf{else}}0.\end{aligned}$
[…]
5.4. Unary value predicate
In B-RASP, the value predicate $V(i,j)$ depends on both $i$ (the query position) and $j$ (the key/value position), but in actual transformers, it depends on $j$ only. As we will see, the dependence on $i$ is convenient for simulating residual connections. But it does not add any expressivity, as can be seen by the fact that the simulation of LTL $\mathrel{\textbf{until}}$ does not use a value predicate that depends on $i$.
More directly, we have the following:
Proposition 5.1.
Any B-RASP program is equivalent to one in which all value predicates $V(i,j)$ depend only on $j$.
Proof.
[…]
The base cases are
$P(i)={\operatorname{{\mathbf{\blacktriangleleft}}}}\left[M(i,j),S(i,j)\right]\;P^{\prime}(j),$
which already has the desired form, and
$P(i)={\operatorname{{\mathbf{\blacktriangleleft}}}}\left[M(i,j),S(i,j)\right]\;P^{\prime}(i),$
which is simply equivalent to $P^{\prime}(i)$. If the value predicate is a conjunction,
$P(i)$ $={\operatorname{{\mathbf{\blacktriangleleft}}}}\left[M(i,j),S(i,j)\right]\;(V_{1}(i,j)\land V_{2}(i,j))$
we may rewrite this as
$P_{1}(i)$ $={\operatorname{{\mathbf{\blacktriangleleft}}}}\left[M(i,j),S(i,j)\right]\;V_{1}(i,j)$
$P_{2}(i)$ $={\operatorname{{\mathbf{\blacktriangleleft}}}}\left[M(i,j),S(i,j)\right]\;V_{2}(i,j)$
$P(i)$ $=P_{1}(i)\land P_{2}(i).$
Similarly for disjunction and negation. ∎
5.5. Non-strict masking
A B-RASP attention operation with non-strict masking has a mask of $i\leq j$ or $j\leq i$. (Consequently, there is always a position that satisfies the mask, so a default position is not needed.)
Unlike the other variants listed above, non-strict masking does reduce the expressivity of B-RASP. First, we note that non-strict masking can be expressed using strict masking:
${\operatorname{{\mathbf{\blacktriangleleft}}}}\left[j\geq i,S(i,j)\right]\;V(i,j)$
${}\equiv{\operatorname{{\mathbf{\blacktriangleleft}}}}\left[j>i,S(i,j)\right]\;(S(i,i)\land V(i,i)\lor\lnot S(i,i)\land V(i,j)).$
We can furthermore show that this inclusion is strict because non-strict masking predicates are severely limited, as suggested by the following.
Proposition 5.2.
The set of vectors $\mathcal{V}=\{Q_{\texttt{EOS}},\lnot Q_{\texttt{EOS}},\mathbf{0},\mathbf{1}\}$ is closed under (a) position-wise operations and (b) attention with non-strict masking.
Proof.
(a) It is easily verified that $\mathcal{V}$ is closed under position-wise operations. (b) In the case of attention operations, we see that either every non-EOS position attends to EOS, or every non-EOS position attends to some non-EOS position. This is because (i) all atomic subformulas of $S(i,j)$ depending on $i$ have identical values for $i\leq n$ and (ii) $M(i,n+1)$ is identical for every $i\leq n$, so if one non-EOS position attends to EOS, all of them must as well.
Now, by Proposition 5.1 we may assume that the value predicate depends only on $j$. Regardless of the choice of $V(j)$, the first $n$ elements of the resulting vector will be identical, so it must be in $\mathcal{V}$. ∎
Intuitively, this means that non-strict masked operations are unable to distinguish between any positions other than EOS. This is not the case with strict masking, so even very simple languages will separate non-strict masking from strict masking.
Theorem 5.3.
B-RASP with strict masking is strictly more expressive than with non-strict masking.
Proof.
No B-RASP program with non-strict masking can recognize the language $\{a\}$ with $\Sigma=\{a\}$. To begin with, we note that $Q_{a}\equiv\lnot Q_{\texttt{EOS}}$, so by Proposition 5.2 all definable vectors are either $Q_{\texttt{EOS}},\lnot Q_{\texttt{EOS}},\mathbf{0}$, or $\mathbf{1}$. This means the program accepts $a\cdot\texttt{EOS}$ iff it accepts $aa\cdot\texttt{EOS}$, as there are no vectors which can distinguish between them. Therefore, no such program can recognize the language $\{a\}$.
Secondly, the following B-RASP program with strict masking accepts $\{a\}$ by accepting strings where the leftmost $a$ is also the $n$-th symbol. We use $Y$ as the output vector:
$P(i)$ $={\operatorname{{\mathbf{\blacktriangleleft}}}}\left[j>i,1\right]\;(Q_{\texttt{EOS}}(j)\land Q_{a}(i))$
$Y(i)$ $={\operatorname{{\mathbf{\blacktriangleleft}}}}\left[1,Q_{a}(j)\right]\;P(j).$
This argument has similarities with Bhattamishra et al. (2020, §7.1), which argues that non-strict masked soft-attention transformers without positional encodings cannot recognize Dyck-1 of depth 2 because the soft-attention layer assigns the same value to every position. Thus, the transformer cannot distinguish between consecutive occurrences of a single symbol at the start of a sentence. However, the addition of a BOS symbol overcomes this limitation, as observed empirically by Ebrahimi et al. (2020). The similarities between how strict masking for hard attention as well as the addition of BOS for soft attention transformers both enable positions to be distinguished is notable for future investigation.
6. Masked Boolean transformers
In this section, we prove that masked Boolean transformers, which have activation values in $\{0,1\}$, are equivalent to B-RASP.
6.1. Definitions
First, we define a more abstract notion of hard-attention transformer which may be specialized to define particular models. An abstract masked hard-attention transformer model is specified by the following components.
  1. (1)
    A set $\mathcal{A}$ of activation values and a set $\mathcal{S}$ of attention score values.
  2. (2)
    A set $\mathcal{F}_{P}$ of position-wise functions mapping $\mathcal{A}$} to $\mathcal{A}$}.
  3. (3)
    A set $\mathcal{F}_{S}$ of score functions mapping $\mathcal{A}\times\mathcal{A}$ to $\mathcal{S}$.
  4. (4)
    A set $\mathcal{F}_{V}$ of value functions mapping ${\mathcal{A}}\times\mathcal{A}$ to $\mathcal{A}$}.
  5. (5)
    A set $\mathcal{M}$ of masks, each of which is a map of $\mathbb{N}_{>0}\times\mathbb{N}_{>0}$ to $\{0,1\}$.
  6. (6)
    A set $\mathcal{F}_{C}$ of choice functions, which map any sequence $(s_{1},\ldots,s_{n+1})$ of values from $\mathcal{S}\cup\{\bot\}$ to a value in $[n+1]$, where $\bot$ is a value not in $\mathcal{S}$.
Note that the only positional information is conveyed by masking.
A layer is a length-preserving function mapping $\mathcal{A}^{*}$ to $\mathcal{A}^{*}$. A position-wise layer is specified by a function $f_{P}\in\mathcal{F}_{P}$, which we can extend to a position-wise function on sequences:
$f_{P}(v_{1},\ldots,v_{n+1})=f_{P}(v_{1}),\ldots,f_{P}(v_{n+1}).$
An attention layer is specified by a mask $M\in\mathcal{M}$, a score function $f_{S}\in\mathcal{F}_{S}$, a choice function $f_{C}\in\mathcal{F}_{C}$, and a value function $f_{V}\in\mathcal{F}_{V}$. On input $v_{1},\ldots,v_{n+1}$, the output values $v_{1}^{\prime},\ldots,v_{n+1}^{\prime}$ for the attention layer are computed as follows. Fix $i\in[n+1]$. For $j\in[n+1]$, if $M(i,j)=1$ let $s_{i,j}=f_{S}(v_{i},v_{j})$, else let $s_{i,j}=\bot$. Apply $f_{C}$ to the sequence $(s_{i,1},\ldots s_{i,n+1})$ to get an index $j_{i}\in[n+1]$. Then the value of $v_{i}^{\prime}$ is $f_{V}(v_{i},v_{j_{i}})$.
Note that this definition has one attention head per attention layer; multi-head attention could be simulated by combining the results of multiple successive attention layers.
An instance of an abstract masked hard-attention transformer model, which we simply call an abstract masked hard-attention transformer, is a length preserving map from $\Sigma_{\texttt{EOS}}^{*}$ to $\mathcal{A}^{*}$ that is the composition of an embedding function and a finite sequence of position-wise layers and attention layers. The embedding function is a function $e\colon\Sigma\to\mathcal{A}$, which is extended to a position-wise function on strings.
A transformer $T$ is run on a string $w$ by applying it to $w\cdot\texttt{EOS}$. We write $T(w)]_{i}$ for the final activation value at position $i$ when $T$ is run on input $w$. To use a transformer as a language recognizer, we need an output function that maps $\mathcal{A}^{*}$ to $\{0,1\}$.
Next, we define our first specialization of the above transformer model. For each positive integer $d$, we define a masked Boolean transformer model of dimension $d$ as follows.
  • The activation values $\mathcal{A}$ are the Boolean vectors $\{0,1\}^{d}$.
  • The score values $\mathcal{S}$ are the Boolean values $\{0,1\}$.
  • The position-wise functions $\mathcal{F}_{P}$ are all functions from $\mathcal{A}$ to $\mathcal{A}$.
  • The score functions $\mathcal{F}_{S}$ are all functions from $\mathcal{A}\times\mathcal{A}$ to $\mathcal{S}$.
  • The value functions $\mathcal{F}_{V}$ are all functions from $\mathcal{A}\times\mathcal{A}$ to $\mathcal{A}$.
  • There are three types of masks, as follows.
    1. (1)
      none: $M(i,j)=1$.
    2. (2)
      future-masking: $M(i,j)=(j<i)$.
    3. (3)
      past-masking: $M(i,j)=(i<j)$.
  • $\mathcal{F}_{C}$ contains two possible choice functions: leftmost maximum and rightmost maximum. Given scores $(s_{i,1},\ldots s_{i,n+1})$, if all the values are $\bot$, then the index $j_{i}$ chosen is $n+1$. Otherwise for leftmost maximum (resp. rightmost maximum), $j_{i}$ is the least (resp., greatest) index $j\in[n+1]$ that maximizes $s_{i,j}$, where the ordering is $\bot<0<1$.
Since the activation values are now vectors, we write $T(w)]_{i,k}$ for the $k$-th coordinate of $T(w)]_{i}$.
6.2. B-RASP to masked Boolean transformers
Let $p$ be a B-RASP program with Boolean vectors $P_{1},\ldots,P_{d}$. We say that a masked Boolean transformer $T$ with $d^{\prime}\geq d$ dimensions simulates $p$ iff for every input $w\in\Sigma^{*}$ with length $n$, we have, for all $i\in[n+1]$ and $k\in[d]$,
$[T(w)]_{i,k}=\begin{cases}1&\text{if $w\models P_{k}(i)$}\\ 0&\text{otherwise.}\end{cases}$
Theorem 6.1.
Let $p$ be a B-RASP program. There exists a masked Boolean transformer $T_{p}$ that simulates $p$.
Proof.
We prove that the first $k$ operations of $p$ can be simulated by a transformer, by induction on $k$. The base case $k=0$ is trivial. Assume that the Boolean vectors $P_{1}(i),\ldots,P_{k}(i)$ are simulated by a masked Boolean transformer. We want to show that $P_{k+1}(i)$ can be simulated as well.
If $P_{k+1}$ is a built-in Boolean vector $Q_{a}$ for some $a\in\Sigma_{\texttt{EOS}}$, then set the embedding $e(a)$ to the one-hot vector with a 1 in the $(k+1)$-st coordinate and a 0 in all other coordinates.
If $P_{k+1}(i)=f(P_{1}(i),\ldots,P_{k}(i))$ where $f$ is some Boolean function, add to $T_{p}$ a position-wise layer that computes the function that performs the same Boolean operation on the first $k$ coordinates and routes the resulting value to the $(k+1)$-st coordinate:
$f^{\prime}\left(\begin{bmatrix}v_{1}\\ \vdots\\ v_{k}\\ 0\\ 0\\ \vdots\end{bmatrix}\right)=\begin{bmatrix}v_{1}\\ \vdots\\ v_{k}\\ f(v_{1},\ldots,v_{k})\\ 0\\ \vdots\end{bmatrix}$.
The last case is if $P_{k+1}$ is an attention operation, that is,
$P_{k+1}(i)$ $={\operatorname{{\mathbf{\blacktriangleleft}}}}\left[M(i,j),S(i,j)\right]\;V(i,j)$
$P_{k+1}(i)$ $={\operatorname{{\mathbf{\blacktriangleright}}}}\left[M(i,j),S(i,j)\right]\;V(i,j).$
This is simulated in $T_{p}$ using an attention layer. The score predicate $S(i,j)$ is a Boolean combination of previous Boolean vectors:
$S(i,j)=g(P_{1}(i),\ldots,P_{k}(i),P_{1}(j),\ldots,P_{k}(j)).$
The score function $f_{S}(v,w)$ computes the same Boolean combination of coordinates in $v$ and $w$:
$f_{S}\left(\begin{bmatrix}v_{1}\\ \vdots\\ v_{k}\\ 0\\ \vdots\end{bmatrix},\begin{bmatrix}w_{1}\\ \vdots\\ w_{k}\\ 0\\ \vdots\end{bmatrix}\right)=g(v_{1},\ldots,v_{k},w_{1},\ldots,w_{k}).$
The value predicate $V(i,j)$ of $p$ is a Boolean combination of previous predicate vectors: $V(i,j)=g(P_{1}(i),\ldots,P_{k}(i),P_{1}(j),\ldots,P_{k}(j))$. The value function $f_{V}(v,w)$ computes the same Boolean combination of values by the method used for the score predicate. This single-bit value is routed to the coordinate corresponding to $P_{k+1}$, and the other coordinates are copied from $v$.
$f_{V}\left(\begin{bmatrix}v_{1}\\ \vdots\\ v_{k}\\ 0\\ 0\\ \vdots\end{bmatrix},\begin{bmatrix}w_{1}\\ \vdots\\ w_{k}\\ 0\\ 0\\ \vdots\end{bmatrix}\right)=\begin{bmatrix}v_{1}\\ \vdots\\ v_{k}\\ g(v_{1},\ldots,v_{k},w_{1},\ldots,w_{k})\\ 0\\ \vdots\end{bmatrix}$.
[…]
After each operation of $p$, the activation values of the constructed sequence of layers of $T_{p}$ represent the activation vectors that have been assigned values by $p$ to that point. After all the operations of $p$, the final activation values of $T_{p}$ represent the output activation vectors $Y_{\gamma}$. ∎
6.3. Masked Boolean transformers to B-RASP
A B-RASP program $p$ simulates a masked Boolean transformer $T$ of dimension $d$ if in $p$ there are Boolean vectors $Y_{k}$ for $k\in[d]$ such that for any input $w\in\Sigma^{*}$ with length $n$, we have, for all $i\in[n+1]$ and $k\in[d]$, $w\models Y_{k}(i)$ iff $T(w)]_{i,k}=1$.
Theorem 6.2.
Let $T$ be a masked Boolean transformer. There exists a B-RASP program $p_{T}$ that simulates $T$.
Proof.
By induction on the number of layers $L$ in $T$. The base case $L=0$ is just the input embedding function $e$, which we simulate with
$E_{k}(i)=\bigwedge_{a\in\Sigma_{\texttt{EOS}}}Q_{a}(i)\rightarrow[e(a)]_{k}.$
Assume $T$ is a masked Boolean transformer of dimension $d$. The program $p_{T}$ has two Boolean vectors $X_{k}$ and $X_{k}^{\prime}$ for each dimension $k\in[d]$ of the activation vectors; call these the primary and secondary Boolean vectors.
Assume that the first $L$ layers of $T$ are simulated by $p_{T}$. We show how to simulate layer $(L+1)$ by adding operations to $p_{T}$.
Consider a position-wise layer of $T$ given by $f_{P}\colon\{0,1\}^{d}\to\{0,1\}^{d}$. Let $f^{\prime}_{k}(x_{1},\ldots,x_{d})$ for $k\in[d]$ be Boolean formulas equivalent to $f_{P}$. Then add the operations
$X^{\prime}_{L+1,k}(i)$ $=f^{\prime}_{k}\left(X^{\prime}_{L,1}(i),\ldots,X^{\prime}_{L,d}(i)\right)$ $k$ $\in[d]$
Consider an attention layer with mask $M(i,j)$, score function $f_{S}(v,w)$, choice of rightmost or leftmost hard attention, and value function $f_{V}\colon\{0,1\}^{d}\times\{0,1\}^{d}\to\{0,1\}^{d}$. In $p_{T}$ we construct an attention operation. If the attention layer uses leftmost-hard attention, then we construct
$X^{\prime}_{L+1,k}(i)$ $={\operatorname{{\mathbf{\blacktriangleleft}}}}\left[M(i,j),S^{\prime}(i,j)\right]\;V^{\prime}_{k}(i,j)$ $k$ $\in[d].$
[…]
7. Masked hard-attention Transformers
7.1. B-RASP to masked hard-attention transformers
We define a more realistic restriction of an abstract masked hard-attention transformer. For each positive integer $d$, we define a masked hard-attention transformer model of dimension $d$ as follows.
  • The activation values $\mathcal{A}$ are $\mathbb{R}^{d}$.
  • The score values $\mathcal{S}$ are $\mathbb{R}$.
  • The position-wise functions $\mathcal{F}_{P}$ are the functions computable by feed forward neural nets of $3$ layers with ReLU activation.
  • The score functions $\mathcal{F}_{S}$ are all the bilinear functions $\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$, that is, functions that can be written in the form $f(u,v)=u^{\top}Sv$ where $S\in\mathbb{R}^{d\times d}$. (In the usual terminology of transformers, $S$ is factored into $S={W^{\text{Q}}}^{\top}W^{\text{K}}$, so that $f(u,v)=(W^{\text{Q}}u)^{\top}(W^{K}\text{v})$, and $W^{\text{Q}}u$ is a query and $W^{\text{K}}v$ is a key.)
  • The value functions $\mathcal{F}_{V}$ are the functions $\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ that can be written in the form $f(u,v)=u+W^{\text{V}}v$ where $W^{\text{V}}\in\mathbb{R}^{d\times d}$. (In the usual terminology of transformers, the $u$ term is the residual connection, and the $W^{\text{V}}v$ term is the value.)
  • The model permits the same three types of masks as masked Boolean transformers.
  • $\mathcal{F}_{C}$ contains the same two choice functions as for masked Boolean transformers: leftmost and rightmost maximum.
We note a few differences from the standard definition (Vaswani et al., 2017). The fact that our feed-forward networks have 3 layers instead of the standard 2 is a minor difference; we can use two 2-layer networks to simulate one 3-layer network. Our definition lacks layer normalization and position embeddings; please see Section 8 for discussion of these differences. Finally, regarding the fact that our attention masking is strict, whereas standard masking is non-strict, the remarks in Section 5.5 apply.
A masked hard-attention transformer simulates a B-RASP program $p$ in exactly the same sense that a masked Boolean transformer does.
Theorem 7.1.
Let $p$ be a B-RASP program. There exists a masked hard-attention transformer $T_{p}$ that simulates $p$.
Proof.
We prove that the first $k$ operations of $p$ can be simulated by a masked hard-attention transformer, by induction on $k$. The base case $k=0$ is trivial. Assume that the Boolean vectors $P_{1},\ldots,P_{k}$ can be simulated by a masked hard-attention transformer. We want to show that $P_{k+1}$ can be simulated as well.
If $P_{k+1}(i)$ is a built-in predicate $Q_{a}$ for some $a\in\Sigma_{\texttt{EOS}}$, then (as in Theorem 6.1) set the embedding $e(a)$ to the one-hot vector with a 1 in the $(k+1)$-st coordinate and a 0 in the other coordinates.
If $P_{k+1}(i)$ is a Boolean combination of $\{P_{1}(i),\ldots,P_{k}(i)\}$, it can be put into DNF and computed by a $3$-layer FFN with ReLU activation functions. The first layer computes all the negations and conjunctions, using the fact that for any Boolean values $b_{1}$, $b_{2},\ldots,b_{m}$ we have
$\neg b_{1}$ $=1-b_{1}$
$b_{1}\wedge b_{2}\wedge\ldots\wedge b_{m}$ $=\operatorname{{\bf ReLU}}(b_{1}+b_{2}+\ldots+b_{m}-(m-1))$
and the second and third layers compute the disjunction, using the fact that
$b_{1}\vee b_{2}\vee\ldots\vee b_{m}$ $=1-\operatorname{{\bf ReLU}}(1-(b_{1}+b_{2}+\ldots+b_{m})).$
If $P_{k+1}$ is an attention operation, either of
$P_{k+1}(i)$ $={\operatorname{{\mathbf{\blacktriangleleft}}}}\left[M(i,j),S(i,j)\right]\;V(i,j)$
$P_{k+1}(i)$ $={\operatorname{{\mathbf{\blacktriangleright}}}}\left[M(i,j),S(i,j)\right]\;V(i,j)$
we construct a leftmost- or rightmost-hard attention layer with mask $M(i,j)$.
The score predicate $S(i,j)$ can be written without loss of generality in canonical DNF form, that is, where each clause is a conjunction of one positive or negative literal for each of the $P_{\ell}(i)$ and $P_{\ell}(j)$, so that at most one clause of the formula is $1$. We separate each clause of $S(i,j)$ into a conjunction of literals depending on $i$ and a conjunction of literals depending on $j$. Thus,
$S(i,j)$ $=\bigvee_{\ell=1}^{m}\left(\alpha_{\ell}(P_{1}(i),\ldots,P_{k}(i))\land\beta_{\ell}(P_{1}(j),\ldots,P_{k}(j))\right).$
The value predicate $V(i,j)$ can be written without loss of generality, by Proposition 5.1, as
$V(i,j)=g(P_{1}(j),\dots,P_{k}(j)).$
Then the position-wise layer uses $2m+1$ dimensions of scratch space to compute the function
[…]
The score function is $f_{S}(u,v)=u^{\top}S^{\prime}v$, where the matrix $S^{\prime}$ is all zero, except that $S^{\prime}_{r,r+1}$ is set to $1$ if for some $\ell$ with $1\leq\ell\leq m$, coordinate $r$ represents $\alpha_{\ell}$ and coordinate $r+1$ represents $\beta_{\ell}$. This ensures that
[…]
$\qquad\sum_{\ell=1}^{m}\alpha_{\ell}(P_{1}(i),\ldots,P_{k}(i))\,\beta_{\ell}(P_{1}(j),\ldots,P_{k}(j))=S(i,j).$
The value function is $f_{V}(u,v)=u+W^{\text{V}}v$ where $W^{\text{V}}]_{k+1,r}=1$ if coordinate $r$ represents $g$ and all other entries are 0. This ensures that
$f_{V}\left(f_{P}\left(\begin{bmatrix}P_{1}(i)\\ \vdots\\ P_{k}(i)\\ 0\\ \vdots\end{bmatrix}\right),f_{P}\left(\begin{bmatrix}P_{1}(j)\\ \vdots\\ P_{k}(j)\\ 0\\ \vdots\end{bmatrix}\right)\right)=\begin{bmatrix}P_{1}(i)\\ \vdots\\ P_{k}(i)\\ V(i,j)\\ \vdots\end{bmatrix}$.
Thus, for any input $w\in\Sigma^{*}$, the final activation vector of $T_{p}$ at position $i$ is equal to the output of $p$ at position $i$. ∎
7.2. Masked hard-attention transformers to B-RASP
To convert a masked hard-attention transformer to B-RASP, we first show that all of the intermediate values computed by the transformer are drawn from a finite set and therefore can be represented using $O(1)$ bits of precision. Hao et al. (2022) previously proved that hard-attention transformers use $O(\log n)$ bits; the difference is that they assumed arbitrary position embeddings, but our transformers do not have position embeddings. (See Section 8.1 below for discussion of adding position embeddings.)
Lemma 7.2.
Let $T$ be a masked hard-attention transformer. There is a finite set $\mathbb{F}\subseteq\mathbb{R}$ such that for all input strings $w$, all the attention scores and components of activation vectors computed by $T$ on input $w$ belong to $\mathbb{F}$.
Proof.
We prove that regardless of the input, layer $L$ has at most $(|\Sigma|+1)^{2^{L}}$ different possible output activation vectors, by induction on the depth of $T$. The base case is just the embedding function. Since there are no position embeddings, the embedding at position $i$ is determined entirely by $w_{i}$, so there are at most $(|\Sigma|+1)$ possible activation vectors.
Assume that $T$ has $(L+1)$ layers and the output of layer $L$ has at most $(|\Sigma|+1)^{2^{L}}$ possible activation vectors. If layer $(L+1)$ is a position-wise layer, its output clearly also has at most $(|\Sigma|+1)^{2^{L}}$ possible activation vectors. If layer $(L+1)$ is a self-attention layer, the output at position $i$ depends only on layer $L$’s outputs at positions $i$ and $j_{i}$, where $j_{i}$ is the position that $i$ attends to. Thus the output has at most $\left((|\Sigma|+1))^{2^{L}}\right)^{2}=(|\Sigma|+1)^{2^{L+1}}$ possible activation vectors.
Similarly, at layer $(L+1)$, every attention score depends on two activation vectors from layer $L$, so there are at most $(|\Sigma|+1)^{2^{L+1}}$ possible scores. Then $\mathbb{F}$ is the union over all layers of the possible attention scores and components of the possible activation vectors. ∎
So any attention score or component of an activation vector can be represented using $\lceil\log_{2}|\mathbb{F}|\rceil$ bits. Define a mapping $\langle\mathord{\cdot}\rangle\colon\mathbb{F}\to\{0,\ldots,2^{B}-1\}$ such that $u<v$ iff $\langle u\rangle<\langle v\rangle$, and write $\langle v\rangle_{b}$ for the bit of $\langle v\rangle$ with place value $2^{b}$.
Let $T$ be a masked hard-attention transformer with $d$ dimensions. Let $B$ be the number of bits needed to store $T$’s activation vector components and attention scores, by Lemma 7.2. A B-RASP program $p$ simulates $T$ if in $p$ there are Boolean vectors $Y_{k,b}$ for $k\in[d]$ and $0\leq b<B$ such that for any input $w\in\Sigma^{*}$ of length $n$, for all $i\in[n+1]$, $k\in[d]$, and $0\leq b<B$, we have $w\models Y_{k,b}(i)$ iff $\langle[T(w)]_{i,k}\rangle_{b}=1$.
Theorem 7.3.
For any masked hard-attention transformer $T$, there is a B-RASP program $p_{T}$ that simulates $T$.
Proof.
We proceed by induction on the depth of $T$. The base case is the input embedding function $e$, which is simulated by predicates
$E_{k,b}(i)=\bigwedge_{a\in\Sigma_{\texttt{EOS}}}Q_{a}(i)\rightarrow\langle[e(a)]_{k}\rangle_{b}.$
Assume that the first $L$ layers of $T$ are simulated by a program $p$. We extend $p$ to simulate layer $(L+1)$ as follows.
If layer $(L+1)$ is a position-wise layer, implementing it in B-RASP is straightforward.
If layer $(L+1)$ is a rightmost-hard attention layer with mask $M(i,j)$, assume that the score function has been converted to Boolean expressions $S^{\prime}_{b}(i,j)$ for the $b$-th bit of the score for positions $i$ and $j$, and the value function has been converted to Boolean expressions $V^{\prime}_{c,b}(i,j)$ for the $b$-bit of the $c$-th coordinate of the value.
We define a sequence of predicates to compute the maximum of $S^{\prime}(i,j)$ for all $j$, and to test whether a particular $j$ is a maximizer:
$\text{Argmax}_{B}(i,j)$ $=1$
For $b=B-1,B-2,\ldots,0$:
$\text{Max}_{b}(i)$ $=\begin{aligned} {\operatorname{{\mathbf{\blacktriangleright}}}}&\left[M(i,j),\text{Argmax}_{b+1}(i,j)\land S^{\prime}_{b}(i,j)\right]\\ &(\text{Argmax}_{b+1}(i,j)\land S^{\prime}_{b}(i,j))\end{aligned}$
$\text{Argmax}_{b}(i,j)$ $=\bigwedge_{b^{\prime}=b}^{B-1}\left(S^{\prime}_{b^{\prime}}(i,j)\leftrightarrow\text{Max}_{b^{\prime}}(i)\right).$
Finally, we define a predicate that simulates attention:
$\text{Att}_{c,b}(i)$ $={\operatorname{{\mathbf{\blacktriangleright}}}}\left[M(i,j),\text{Argmax}_{0}(i,j)\right]\;V^{\prime}_{c,b}(i,j).$
[…]
8. Transformer Variants
8.1. Position embeddings
Our definition of transformer does not, so far, include position embeddings; all information about ordering comes from attention masking. A position embedding is a family of functions $\Theta=(\theta_{n})_{n\geq 0}$ where $\theta_{n}(i)$ is a scalar or vector representation of position $i$ in a string of length $n$. We say that $\Theta$ has finite image if $\bigcup_{n\geq 0}\mathop{\textrm{Im}}\theta_{n}$ is finite. In general, our results extend to transformers with any position embedding that has finite image. The class of languages recognized may grow, and we give a recipe for characterizing the new class of languages.
In B-RASP and $\textbf{FO}[\mathord{<}]$, we can add built-in predicates as follows. Let $\Pi=(\pi_{n})_{n\geq 0}$ be a family of functions $\pi_{n}\colon[n+1]\to\{0,1\}$. In both $\textbf{B-RASP}[\Pi]$ and $\textbf{FO}[\mathord{<},\Pi]$, there is an additional predicate symbol $\Pi$ such that $w\models\Pi(i)$ iff $\pi_{n}(i)=1$, where $n=|w|$.
For example, if
$\text{Mid}_{n}(i)$ $=\begin{cases}1&\text{if $n$ is even and $i=1+n/2$}\\ 0&\text{otherwise}\end{cases}$
then we can define the language $\{a^{m}b^{m}\mid m\geq 0\}$ in $\textbf{FO}[\mathord{<},\text{Mid}]$ as:
$(\exists i)(\text{Mid}(i)\land(\forall j)\begin{aligned} &((j<i\rightarrow\neg Q_{b}(j))\\ &{}\land(j>i\lor j=i\rightarrow\neg Q_{a}(j)))).\end{aligned}$
A similar definition could be written in $\textbf{B-RASP}[\text{Mid}]$.
Theorem 8.1.
Let $\Theta=(\theta_{n})_{n\geq 0}$ be a position embedding with finite image. There exists a collection of predicates $\mathcal{P}$ such that the following classes of languages are the same:
  • languages recognized by masked hard-attention transformers with position embedding $\Theta$
  • languages defined by $\textbf{B-RASP}[\mathcal{P}]$
  • languages defined by $\textbf{FO}[\mathord{<},\mathcal{P}]$
Proof.
Because $\Theta$ has finite image, Lemma 7.2 still holds for any masked hard-attention transformer with position embedding $\Theta$. Let $\mathcal{P}$ be the collection of predicates that test whether the $b$-th bit of the $k$-th coordinate of $\theta_{n}(i)$ is set. The proof of equivalence of masked hard-attention transformers with B-RASP extends easily to equivalence of masked hard-attention transformers with position embedding $\Theta$ and $\textbf{B-RASP}[\mathcal{P}]$. When converting a transformer to a $\textbf{B-RASP}[\mathcal{P}]$ program, we represent each coordinate of $\Theta$ with $B$ predicates from $\mathcal{P}$. When converting a $\textbf{B-RASP}[\mathcal{P}]$ program to a transformer, we represent each predicate in $\mathcal{P}$ with its own coordinate, whose value is in $\{0,1\}$.
Since Theorem 4.1, Theorem 4.3 and Kamp’s Theorem (Kamp, 1968) hold for any collection of unary predicate symbols, $\textbf{B-RASP}[\mathcal{P}]$ is equivalent to $\textbf{FO}[\mathord{<},\mathcal{P}]$. ∎
Next, we discuss some particular position embeddings.
Sinusoidal position embeddings
The original transformer (Vaswani et al., 2017) used position embeddings with coordinates of the form $\sin(2\pi fi)$ or $\cos(2\pi fi)$. If the $f$’s are rational (and, admittedly, in the original definition, they were not), then the position embeddings form a finite set, so Lemma 7.2 still holds. For any even $d$, let us define a sinusoidal positional embedding with $d$ dimensions to be a position embedding $\Theta=(\theta_{n})_{n\geq 0}$ where
$\theta_{n}(i)=\begin{bmatrix}\sin 2\pi f_{1}\\ \cos 2\pi f_{1}\\ \vdots\\ \sin 2\pi f_{d/2}\\ \cos 2\pi f_{d/2}\end{bmatrix}$
and all the $f_{c}$ for $c\in[d/2]$ are rational.
Let MOD be the collection of predicates $\text{MOD}^{r}_{m}(i)$ for all $0\leq r<m$, which hold just in case $i\equiv r\pmod{m}$.
Corollary 8.2.
Masked hard-attention transformers with sinusoidal position embeddings recognize exactly the regular languages in $\textbf{AC}^{0}$.
Proof.
Let $\Theta$ be a sinusoidal positional embedding. Since the $f_{c}$ are rational, $\Theta$ has finite image. By Theorem 8.1, transformers with positional embedding $\Theta$ are equivalent to $\textbf{FO}[\mathord{<},\mathcal{P}]$.
It’s easy to see that every predicate in $\mathcal{P}$ can be expressed in terms of MOD; for the converse, observe that we can use a 2-layer ReLU network to compute $\text{MOD}^{r}_{m}$ (Chiang et al., 2023, Lemma 20):
$h(i)$ $=\operatorname{{\bf ReLU}}\left(\sin\frac{2\pi r}{m}\sin\frac{2\pi i}{m}+\cos\frac{2\pi r}{m}\cos\frac{2\pi i}{m}-\cos\frac{2\pi}{m}\right)$
$=\operatorname{{\bf ReLU}}(\cos(2\pi(i-r)/m))$
$\text{MOD}^{r}_{m}(i)$ $=\left(1-\cos\frac{2\pi}{m}\right)h(i).$
Thus transformers with sinusoidal positional embeddings are equivalent to $\textbf{FO}[\mathord{<},\text{MOD}]$. And $\textbf{FO}[\mathord{<},\text{MOD}]$, in turn, defines exactly the class of regular languages in $\textbf{AC}^{0}$ (Mix Barrington et al., 1992). ∎
Arbitrary position embeddings
Finally, we may consider arbitrary position embeddings, subject to the condition of finite image. The corresponding collection of predicates is the set of all possible monadic predicates, which we call Mon. [^2] Although Mix Barrington et al. (2005) define Mon to be the collection of all monadic predicates without dependence on $n$, Barceló et al. (2024) do allow monadic predicates to depend on $n$.
Corollary 8.3.
Masked hard-attention transformers that have position embeddings with finite image recognize exactly the languages definable in $\textbf{FO}[\mathord{<},\text{Mon}]$.
Barceló et al. (2024) show that any language definable in $\textbf{FO}[\mathord{<},\text{Mon}]$ can be recognized by a hard-attention transformer without attention masking and with some position embedding, but left the other direction as an open question. Here, by making use of attention masking and restricting position embeddings to those with finite image, we have obtained an exact characterization.
Indeed, either infinite image or attention masking is needed in order to recognize every language in $\textbf{FO}[\mathord{<},\text{Mon}]$. Without either, there must be two positions $i$ and $j$ with the same position embedding, by the pigeonhole principle, so unmasked attention would not be able to distinguish one string with $a$ and $b$ at positions $i$ and $j$ and another string with $a$ and $b$ at positions $j$ and $i$. So no masked hard-attention transformer with position embeddings with finite image and unmasked attention can recognize the language $a^{*}b^{*}$, but this language is definable even in $\textbf{FO}[\mathord{<}]$.
8.2. Layer normalization
Standard transformers (Vaswani et al., 2017) include layer normalization (Ba et al., 2016), but our definition above does not. Since layer normalization is a position-wise function, the proof of Theorem 7.3 is unaffected. But the construction of Theorem 7.1 does need to be modified to circumvent layer normalization (cf. Chiang et al., 2023, Proposition 22). Previously, we used 1 to represent true and 0 to represent false; now, we use a pair of activations to represent a truth value, $(1,0)$ for true and $(0,1)$ for false. This ensures that every vector has mean and variance independent of the input $w$, so we can set the parameters of each layer normalization so that it has no effect.
9. Counter-Free Automata to B-RASP
Finally, we give an alternative proof that any star-free language is definable in B-RASP (Corollary 4.4), using the fact that the star-free languages are exactly those recognized by counter-free automata, and a counter-free automaton can be decomposed using Krohn-Rhodes theory into a cascade of identity-reset automata. We will show that such a cascade can be simulated by a B-RASP program.
A deterministic finite automaton (DFA) is a tuple $A=(\Sigma,Q,\delta)$, where $\Sigma$ is the input alphabet, $Q$ is the finite set of states, and $\delta\colon Q\times\Sigma\rightarrow Q$ is the transition function. A counter-free automaton is a DFA in which no string induces a permutation on $Q$ other than the identity. Schützenberger (1965) proved that the star-free languages are exactly those recognized by counter-free automata.
Maler (2010) gives the following automata-theoretic version of the Krohn-Rhodes decomposition theorem, which we explain below.
Theorem 9.1 ((Maler, 2010), Theorem 3).
For every [deterministic finite] automaton $A$ there exists a cascade decomposition
$C=B_{1}\circ B_{2}\circ\cdots\circ B_{k}$
such that the following are true.
  1. (1)
    Each $B_{i}$ is a permutation-reset automaton.
  2. (2)
    There is a homomorphism $\phi$ from $C$ to $A$.
  3. (3)
    Any permutation group in some $B_{i}$ is homomorphic to a subgroup of the transformation semigroup of $A$.
The pair $(C,\phi)$ is called a cascade decomposition of $A$.
[…]
Property (2) says that there is a mapping $\phi$ from the states of $C$ to the states of $A$ such that for any state $q$ of $C$, we have $\phi(\delta_{C}(q,a))=\delta_{A}(\phi(q),a)$.
Property (3) implies that if the automaton $A$ is counter-free, then all of the automata $B_{i}$ in the cascade decomposition are identity-reset automata. An identity-reset automaton is a DFA $A=(\Sigma,Q,\delta)$ such that for every $a\in\Sigma$, the mapping $q\mapsto\delta(q,a)$ is either the identity ($\delta(q,\sigma)=q$ for all $q$) or constant (there is a $q_{a}$ such that $\delta(q,a)=q_{a}$ for all $q$).
For example, consider the automaton $A_{3}$ shown in Figure 6(a). A decomposition of this automaton into a cascade product of three identity-reset automata is shown in Figure 6(b). The global automaton derived from the decomposition is shown in Figure 6(c) with the homomorphism $\phi$ to states of $A_{3}$.
0112233RRLLRRLLRRLLLLRR
(a) Automaton $A_{3}$: move right ($R$), move left ($L$).
AABBCCDDEEFFRRLLRRLLB,RB,RA,LA,LA,LA,LB,RB,RBD,RBD,RAC,LAC,LAC,LAC,LBD,RBD,R
(b) Cascade of identity-reset automata for $A_{3}$. Omitted transitions are self-loops, eg, inputs $A,R$ and $B,L$ are self-loops on $C$ and $D$.
ACE(0)ACE(0)BCE(1)BCE(1)ADE(1)ADE(1)BDE(2)BDE(2)ACF(1)ACF(1)BCF(2)BCF(2)ADF(2)ADF(2)BDF(3)BDF(3)RRLLLLRRLLRRLLRRRRLLLLRRLLRRLLRR
(c) The global automaton of the cascade product in part (b). For state $q$, the number in parentheses is the state $\phi(q)$ in $A_{3}$.
Figure 6. Example automaton and its cascade decomposition.
Let $A=(\Sigma,Q,\delta)$ be a DFA and $s\in Q$. For a word $w_{1}\cdots w_{n}\in\Sigma^{*}$, the sequence of states traversed by $A$ from state $s$ on this input is $q_{0},\ldots,q_{n}$, where $q_{0}=s$ and for each $k$, $q_{k+1}=\delta(q_{k},w_{k})$. A B-RASP program $p$ simulates $A$ started in state $s$ iff for every input word $w\in\Sigma^{*}$, the output Boolean vectors of $p$ on input $w$ encode the sequence of states traversed by $A$ from state $s$ on input $w$. The state at position $i$ is the state before the symbol at position $i$ is read.
Lemma 9.2.
Let $B=(\Sigma,Q,\delta)$ be any identity-reset automaton, and let $s\in Q$ be a start state. There exists a B-RASP program $p_{B}$ that simulates $B$ started in state $s$.
Proof.
For each state $q\in Q$, let $I_{q}$ denote the set of symbols $a$ such that there is a transition into $q$ on $a$ from a different state, and let $O_{q}$ denote the set of symbols $a$ such that there is a transition out of $q$ on $a$ to a different state. Then $I_{q}\cap O_{q}=\emptyset$ because of the identity-reset assumption.
To determine if $B$ is in state $q\neq s$ before reading the input symbol in position $i$, it is sufficient to attend to the closest position $j<i$ that contains a symbol from $I_{q}$ or $O_{q}$, if any. If there are none, then $B$ is not in state $q$ at position $i$, because no transition entered state $q$. If the symbol at position $j$ is from $I_{q}$, then $B$ is in state $q$ at position $i$, because state $q$ was entered and has not since been left. If the symbol at position $j$ is from $O_{q}$, then $B$ is not in state $q$ at position $i$, because state $q$ was left and has not since been re-entered.
The case of state $s$ is slightly different. Attending to the closest position $j<i$ that contains a symbol from $I_{s}$ or $O_{s}$, if the symbol at $j$ is not from $O_{s}$, then $B$ is in state $s$ at position $i$ having either never left $s$ or entered $s$ more recently than it left $s$.
In B-RASP, we can define a Boolean vector $B_{q}(i)$, which is true iff $B$ is in state $q$ at position $i$:
$B_{q}(i)$ $={\operatorname{{\mathbf{\blacktriangleright}}}}\left[j<i,\;\bigvee_{\mathclap{a\in I_{q}\cup O_{q}}}Q_{a}(j)\right]\;\bigvee_{a\in I_{q}}Q_{a}(j)\quad\text{for }q\neq s$
$B_{s}(i)$ $={\operatorname{{\mathbf{\blacktriangleright}}}}\left[j<i,\;\bigvee_{\mathclap{a\in I_{s}\cup O_{s}}}Q_{a}(j)\right]\;\neg\bigvee_{\mathclap{a\in O_{s}}}Q_{a}(j).$
Lemma 9.3.
Let $B_{1}=(\Sigma,Q_{1},\delta_{1})$ be a DFA that can be simulated from state $s_{1}$ by a B-RASP program $p_{B_{1}}$. Let $B_{2}=(Q_{1}\times\Sigma,Q_{2},\delta_{2})$ be an identity-reset automaton and let $C=B_{1}\circ B_{2}$. Then there is a B-RASP program $p_{C}$ that simulates $C$ started in state $(s_{1},s_{2})$ for an arbitrary $s_{2}\in Q_{2}$.
Proof.
Let $B_{1,q}(i)$ be predicates that test whether $B_{1}$ is in state $q$ at position $i$, and let $B_{2,q}(i)$ be predicates that test whether $B_{2}$ is in state $q$ at position $i$ (by Lemma 9.2). Define
$Q^{\prime}_{(q,a)}(i)$ $=B_{1,q}(i)\land Q_{a}(i)$ $q$ $\in Q_{1},a\in\Sigma$
and for every line in the definition of the $B_{2,q}$, replace every occurrence of $Q_{(q,a)}$ with $Q^{\prime}_{(q,a)}$. This yields the definition of new predicates $B^{\prime}_{2,q}$. Then we can define predicates $C_{(q,r)}(i)$ that test whether $C=B_{1}\circ B_{2}$ is in state $(q,r)$ at position $i$:
$C_{(q,r)}(i)$ $=B_{1,q}(i)\land B^{\prime}_{2,r}(i)$ $q$ $\in Q_{1},r\in Q_{2}.$
By induction on $k$ we have the following.
Lemma 9.4.
Let $C=B_{1}\circ B_{2}\circ\cdots\circ B_{k}$ be a cascade product such that each $B_{i}$ is an identity-reset automaton, and $s_{i}$ is a state of $B_{i}$. Then there is a B-RASP program $p_{C}$ that can simulate $C$ from state $(s_{1},s_{2},\ldots,s_{k})$.
If we add instructions to the program in Lemma 9.4 that compute the homomorphism $\phi$ (from property (2) of Theorem 9.1) from the states of the cascade product $C$ to the automaton $A$:
$A_{r}(i)=\bigvee_{\mathclap{\begin{subarray}{c}q\\ \phi(q)=r\end{subarray}}}C_{q}(i)$
then we have the following, which implies Corollary 4.4:
Theorem 9.5.
Let $A$ be a counter-free automaton with start state $s$. Then there is a B-RASP program $p_{A}$ that simulates $A$ started in state $s$.